Suppose u(x1, x2 ) = x1^ax2^1-a

  1. (a) Find the optimal bundle x(p, w) and the indirect utility function v(p, w).

  2. (b) Find the Hicksian demand function h(p, u) and the expenditure function e(p, u).

  3. (c) For the remainder of the problem, suppose α = 4 and w = 5. If p = (2,1), what is5

    the optimal bundle? What is the utility of that bundle? [Leave your answer in terms of fractions and exponents]

  4. (d) Suppose the price of good 2 increases, so that the new price vector is p′ = (2, 2).

    (i) Find the new bundle. What is the utility of that bundle? (ii) Compute the compensating variation of the price change.

    (iii) Compute the equivalent variation of the price change.
Answer #1

TILITY FUNCTION OF CONSUMER UCxy X,) = x,x, WE BUOGETENCOME OF CONSUMER P = PRICE OF GOOD X P2PRICE OF GOOD X2 OPTIMAL BUNDLE*in Pyeongchange Company **CP,P3, W)= y sobie XT (P, P2, W) = 4W X7 LP, P2, WP_ (1-2) W ARE P2 MARSHALLIAN DEMAND FUNCTIONS ..:-) ( het moet u doente => x= fu lands) =) X, CP, P2, 0)= v* o * =) X,CP, 82, UDa bete (18+){ «Prado L HILISIAN DEMAND FUNCTEXPENDITURE UTILITY OF C CONSUMED NDITURE FONCTION DEPENDS ON Y OF CONSUMER AND PRICE OF GOODS ECP., P2, 0) = Pix, +2x2 E(Pi,U(X,,X2) = 2X, MX** UC2, 1) = 232 (28(00:2=1.741 (d) PRICE OF GOOD ci) I ICE OF GOOD X₂ INCREASES TO P2 P=2 P=2 W=5 TOPTIMALNEW PRICES P=2 P₂ =2 CP) U=OLD UTILITY= 1.741 = NEW UTILITY= 1514 OLD PRICE P=2 Pz=1CP) - COMPENSATING VARIATION, CU = ECPLU